# Networked Communication for Mean-Field Games with Function Approximation and Empirical Mean-Field Estimation

### v1

@@ 3.1 Q-network and update
For $cl<0$, $[\cdot]^{0}_{cl}$ is a clipping function used in Munchausen RL to prevent numerical issues if the policy is too close to deterministic, as the log-policy term is otherwise unbounded (Vieillard, Pietquin, and Geist 2020; Wu et al. 2024).
@@ 3.2 Communication and adoption of parameters
We extend the work in Benjamin and Abate (2024) to the function-approximation case, where in our work agents broadcast the parameters of the Q-network that defines their policy, rather than the Q-function table. At each iteration $k$, after independently updating their Q-network and policy (Lines 3-14), agents approximately evaluate their new policies by collecting rewards for $E$ steps, and assign the estimated value to $\sigma^{i}_{k+1}$ (Lines 15-20). They then broadcast their Q-network parameters along with $\sigma^{i}_{k+1}$ (Line 22). Receiving these from their neighbours on the communication network, agents select which set of parameters to adopt by taking a softmax over their own and the received estimate values $\sigma^{j}_{k+1}$ (Lines 23-25). They repeat this broadcast and adoption process for $C_{p}$ rounds, which is distinct from the $C_{e}$ communication rounds discussed in Sec. 4. By this process, decentralised agents can adopt policy parameters that are estimated to be better performing than their own, bringing the benefits to convergence speed discussed in Sec. 5.1.
@@ 4 Mean-field estimation and communication
We first describe the most general version of our algorithm for decentralised estimation of the empirical categorical mean-field distribution, assuming the more general setting where $\mathcal{G}^{obs}_{t}$ applies (see discussion in Sec. 2.3). We subsequently detail how the algorithm can be made more efficient in environments where the more abstract visibility graph $\mathcal{G}^{vis}_{t}$ applies, as in our experimental settings. In both cases, the algorithm runs to generate the observation object when a step is taken in the main Alg. 1, i.e. to produce $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$ for the steps $a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t})$ in Lines 5, 17 and 26. Both versions of the algorithm are subject to implicit assumptions, which we highlight and discuss methods for addressing in our ‘Future Work’ section in Appx. C.
Algorithm for the general setting
Algorithm 2 Mean-field estimation and communication algorithm in general settings
@@
12:  return  mean-field estimates $\forall i$ $\tilde{\hat{\mu}}^{i}_{t}$
@@
The method for the general setting functions is given in Alg 2. In this setting, we assume that each agent is associated with a unique ID to avoid the same agents being counted multiple times. Each agent maintains a ‘count’ vector ${\hat{\upsilon}}^{i}_{t}$ of length $|\mathcal{S}|$ i.e. it is the same shape as the vector denoting the true empirical categorical distribution of agents. Each state position in the vector can hold a list of IDs. Before any actions are taken at each time step $t$, each agent’s count vector ${\hat{\upsilon}}^{i}_{t}$ is initialised as full of $\emptyset$ (‘no count’) markers for each state. Then, for each agent $j$ with which agent $i$ is connected via the observation graph, $i$ places $j$’s unique ID in its count vector in the correct state position. Next, for $C_{e}\geq 0$ communication rounds, agents exchange their local counts with neighbours on the communication network, and merge these counts with their own count vector, filtering out the unique IDs of those that have already been counted. If $C_{e}=0$ then the local count will remain purely independent. By exchanging these partially filled vectors, agents are able to improve their local counts by adding the states of agents that they have not been able to observe directly themselves.
@@
Algorithm 3 Mean-field estimation and communication algorithm for environments with $\mathcal{G}^{vis}_{t}$
@@
13:  return  mean-field estimates $\forall i$ $\tilde{\hat{\mu}}^{i}_{t}$
After the $C_{e}$ communication rounds, each state position ${\hat{\upsilon}}^{i}_{t}[s]$ either still maintains the $\emptyset$ marker if no agents have been counted in this state, or it contains $x_{s}>0$ unique IDs. The local mean-field estimate $\tilde{\hat{\mu}}^{i}_{t}$ is then obtained from ${\hat{\upsilon}}^{i}_{t}$ as follows. All states that have a count $x_{s}$ have this count converted into the categorical probability $x_{s}/N$ (we assume that agents know the total number of agents in the finite population, even if they cannot observe them all at each $t$). The total number of agents counted in ${\hat{\upsilon}}^{i}_{t}$ is given by $counted\_agents$ = $\sum_{s\in\mathcal{S}}x_{s}$, and the agents that have not been observed are $uncounted\_agents$ = $N$ $counted\_agents$. In this general setting, the unobserved agents are assumed to be uniformly distributed across all the states, so $uncounted\_agents/(N\times|\mathcal{S}|)$ is added to all the values in $\tilde{\hat{\mu}}^{i}_{t}$, replacing the $\emptyset$ marker for states for which no agents have been observed.
Algorithm for visibility-based environments
We now describe the differences in our estimation algorithm (Alg. 3) for the subclass of environments where $\mathcal{G}^{vis}_{t}$ applies in place of $\mathcal{G}^{obs}_{t}$, i.e. the mutual observability of agents depends in turn on the mutual visibility of states. Recall the benefit of $\mathcal{G}^{vis}_{t}$ over $\mathcal{G}^{obs}_{t}$ is that the former allows an agent in state $s$ to obtain a correct, complete count $x_{s^{\prime}}\geq 0$ of all the agents in state $s^{\prime}$, for any state $s^{\prime}$ that is visible to $s$ (note the count may be zero). Unique IDs are thus not required as there is no risk of counting the same agent twice when receiving communicated counts: either all agents in $s^{\prime}$ have been counted, or no count has yet been obtained for $s^{\prime}$. This simplifies the algorithm and helps preserve agent anonymity and privacy.
Secondly, uncounted agents cannot be in states for which a count has already been obtained, since the count is complete and correct, even if the count is $x_{s^{\prime}}=0$. Therefore after the $C_{e}$ communication rounds, the uncounted agents $uncounted/N$ need be uniformly distributed only across the positions in the vector that still have the $\emptyset$ marker, and not across all states as in the general setting. This makes the estimation more accurate in this special setting.
5 Experiments
Figure 2: ‘Cluster’ task, population-independent policies, 100x100 grid.
We provide two sets of experiments. The first set showcases that our function-approximation algorithm (Alg. 1) can scale to large state spaces for population-independent policies, and that in such settings networked, communicating agents can outperform purely-independent and even centralised agents, and do so by an even greater margin than in the tabular settings from Benjamin and Abate (2024). The second set demonstrates that Alg. 1 can handle population-dependent policies, as well as the ability of Alg. 3 to practically estimate the mean-field distribution locally.
For the types of game used in our demonstrations we follow the gold standard in prior works on MFGs, namely grid-world environments where agents can move in the four cardinal directions or remain in place (uz Zaman et al. 2023; Lauriere et al. 2022; Algumaei et al. 2023; Laurière 2021; Cui, Fabian, and Koeppl 2023; Lauriere et al. 2022; Benjamin and Abate 2024; Wu et al. 2024). We present results from four coordination tasks defined by the reward/transition functions of the agents - see Appx. A.1 for a full technical description. The first two tasks are those used with population-independent policies in Benjamin and Abate (2024), though while they show results for an 8x8 and a ‘larger’ 16x16 grid, our results for these tasks are for 100x100 and 50x50 grids:
  • Cluster. Agents are rewarded for gathering together. The agents are given no indication where they should cluster, agreeing this themselves over time.
  • Target agreement. Agents are rewarded for visiting any of a given number of targets, but their reward is proportional to the number of other agents co-located at the target. Agents must thus coordinate on which single target they will all meet at to maximise their individual rewards.
We also showcase the ability of our algorithm to handle two more complex tasks, using population-dependent policies and estimated mean-field observations:
  • Evade shark in shoal. At each $t$, a ‘shark’ in the environment takes a step towards the grid point containing the most agents according to the empirical mean-field distribution. The shark’s position forms part of agents’ local states in addition to their own position. Agents are rewarded more for being further from the shark, and also for clustering with other agents. As well as featuring a non-stationary distribution, we add ‘common noise’ to the environment, with the shark taking a random step with probability 0.01. Such noise that affects the local states of all agents in the same way, making the evolution of the distribution stochastic, makes population-independent policies sub-optimal (Laurière et al. 2022).
  • Push object to edge. Agents are rewarded for how close they are to an ‘object’ in the environment, and for how close this object is to the edge of the grid. The object’s position forms part of agents’ local states in addition to their own position. The object moves in a direction with a probability proportional to the number of agents on its opposite side, i.e. agents must coordinate on which side of the object from which to ‘push’ it, to ensure it moves toward the edge of the grid.
In these spatial environments, both the communication network $\mathcal{G}^{comm}_{t}$ and the visibility graph $\mathcal{G}^{vis}_{t}$ are determined by the physical distance from agent $i$; we show plots for various radii, expressed as fractions of the maximum possible distance (the diagonal length of the grid). To demonstrate the benefits of the networked architecture by comparison, we also plot the results of modified versions of our algorithm for centralised and independent learners. In the centralised setting, the Q-network updates of agent $i=1$ are automatically pushed to all other agents, and the true global mean-field distribution is always used in place of the local estimate i.e. $\tilde{\hat{\mu}}^{i}_{t}$ = ${\hat{\mu}}_{t}$. In the independent case, there are no links in $\mathcal{G}^{comm}_{t}$ or $\mathcal{G}^{vis}_{t}$, i.e. $\mathcal{E}^{comm}_{t}=\mathcal{E}^{vis}_{t}=\emptyset$. Hyperparameters are discussed in Appx. A.3.
5.1 Results
We evaluate our experiments via two metrics. Exploitability is the most common metric in works on MFGs, and is a measure of proximity to the MFNE. It quantifies how much a best-responding agent can benefit by deviating from the set of policies that generate the current mean-field distribution, with a decreasing exploitability meaning the population is closer to the MFNE. However, there are several issues with this metric in our setting, such that it may give limited or noisy information (see Appx. A.2 for a full discussion). We thus also give a second metric, as in Benjamin and Abate (2024): the population’s average discounted return. This allows us to compare how much agents are learning policies that increase their returns, even when the exploitability metric gives us limited ability to distinguish between the desirability of the MFNEs to which populations are converging.
Population-independent policies in large state-spaces
In Fig. 1 the networked agents generally have lower exploitability than both centralised and independent agents, while in Fig. 2 the networked agents’ exploitability is similar to that of the other cases. However, the networked agents significantly outperform the other architectures in terms of average return on both tasks, indicating that the communication scheme helps agents to find substantially ‘preferable’ equilibria. Moreover, the margin by which the networked agents can outperform the centralised agents is much greater than in Benjamin and Abate (2024), showing that the benefits of the communication scheme are even greater in non-tabular settings. See Appx. A for further experiments with these large state spaces.
Population-dependent policies in complex environments
Figs. 3 and 4, where agents estimate the mean-field distribution via Alg. 3, differ minimally from Figs. 5 and 6 in Appx. A, where agents directly receive the global mean-field distribution. This shows that our estimation algorithm allows agents to appropriately estimate the distribution, even with only one round of communication for agents to help each other improve their local counts. Only in the ‘push object’ task in Fig. 3, and there only with the smaller broadcast radii, do agents slightly underperform the returns of agents in the global observability case in Fig. 5, as is reasonable.
For the reasons given in Appx. A.2, the exploitability metric gives limited information in the ‘push object’ task in Fig. 3. In the ‘evade’ task in Fig. 4, exploitability suggests that centralised learners outperform the other cases. However, all of the networked cases significantly outperform the independent learners in terms of the average return to which they converge in both tasks. In the ‘push object’ task networked learners also significantly outperform centralised learners in all but the case with the smallest broad communication radius, while in the ‘evade’ task all networked cases perform similarly to the centralised case. Recall though that in the real world a centralised architecture is a strong assumption, a computational bottleneck and single point of failure.
Appendix A Experiments
For reproducibility, the code to run our experiments is provided with our Supplementary Materials, and will be made publicly available upon publication.

### v2

@@ 3.2 Communication and adoption of parameters
We adapt the work in Benjamin and Abate (2023) for the function-approximation case, where in our work agents broadcast the parameters of the Q-network that defines their policy, rather than the Q-function table. At each iteration $k$, after independently updating their Q-network and policy (Lines 3-14), agents approximate the infinite discounted return (Def. 3) of their new policies by collecting rewards for $E$ steps, and assign the finite-step discounted sum to $\sigma^{i}_{k+1}$ (Lines 15-20). They then broadcast their Q-network parameters along with $\sigma^{i}_{k+1}$ (Line 22). Receiving these from neighbours on the network, agents select which set of parameters to adopt by taking a softmax over their own and the received estimate values $\sigma^{j}_{k+1}$ (Lines 23-25). They repeat the process for $C_{p}$ rounds. This allows decentralised agents to adopt policy parameters estimated to perform better than their own, accelerating learning as shown in Sec. 5.
@@ 4 Mean-field estimation and communication
We first describe the most general version of our algorithm for decentralised estimation of the empirical categorical mean-field distribution, assuming the more general setting where $\mathcal{G}^{obs}_{t}$ applies (see discussion in Sec. 2.3). We subsequently detail how the algorithm can be made more efficient in environments where the more abstract visibility graph $\mathcal{G}^{vis}_{t}$ applies, as in our experimental settings. In both cases, the algorithm runs to generate the observation object when a step is taken in the main Alg. 1, i.e. to produce $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$ for the steps $a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t})$ in Lines 5, 17 and 26. Both versions of the algorithm are subject to implicit assumptions, which we highlight and discuss methods for addressing in Sup. Mat. E.
4.0.1 Algorithm for the general setting
Algorithm 2 Mean-field estimation and communication in general settings
@@ 4.0.1 Algorithm for the general setting
12:  return  mean-field estimates $\{\tilde{\hat{\mu}}^{i}_{t}\}_{i=1}^{N}$
@@
In this setting, our method (Alg. 2) assumes each agent is associated with a unique ID to avoid the same agents being counted multiple times. Each agent maintains a ‘count’ vector ${\hat{\upsilon}}^{i}_{t}$ of length $|\mathcal{S}|$ i.e. of the same shape as the vector denoting the true empirical categorical distribution of agents. Each state position in the vector can hold a list of IDs. Before any actions are taken at each time step $t$, each agent’s count vector ${\hat{\upsilon}}^{i}_{t}$ is initialised as full of $\emptyset$ (‘no count’) markers for each state. Then, for each agent $j$ with which agent $i$ is connected via the observation graph, $i$ places $j$’s unique ID in its count vector in the correct state position. Next, for $C_{e}\geq 0$ communication rounds, agents exchange their local counts with neighbours on the communication network, and merge these counts with their own count vector, filtering out the unique IDs of those that have already been counted. If $C_{e}=0$ then the local count will remain purely independent. By exchanging these partially filled vectors, agents are able to improve their local counts by adding the states of agents that they have not been able to observe directly themselves.
After the $C_{e}$ communication rounds, each state position ${\hat{\upsilon}}^{i}_{t}[s]$ either still maintains the $\emptyset$ marker if no agents have been counted in this state, or contains $x_{s}>0$ unique IDs. The local mean-field estimate $\tilde{\hat{\mu}}^{i}_{t}$ is then obtained from ${\hat{\upsilon}}^{i}_{t}$ as follows. All states that have a count $x_{s}$ have this count converted into the categorical probability $x_{s}/N$ (we assume that agents know the total number of agents in the finite population, even if they cannot observe them all at each $t$). The total number of agents counted in ${\hat{\upsilon}}^{i}_{t}$ is given by $counted\_agents$ = $\sum_{s\in\mathcal{S}}x_{s}$, and the agents that have not been observed are $uncounted\_agents$ = $N$ $counted\_agents$. In this general setting, the unobserved agents are assumed to be uniformly distributed across all the states, so $uncounted\_agents/(N\times|\mathcal{S}|)$ is added to all the values in $\tilde{\hat{\mu}}^{i}_{t}$, replacing the $\emptyset$ marker for states for which no agents have been observed.
4.0.2 Algorithm for visibility-based environments
We give now the differences in our estimation algorithm (Alg. 3, Sup. Mat. A) for the subclass of environments where $\mathcal{G}^{vis}_{t}$ applies in place of $\mathcal{G}^{obs}_{t}$, i.e. the mutual observability of agents depends in turn on the mutual visibility of states. The benefit of $\mathcal{G}^{vis}_{t}$ over $\mathcal{G}^{obs}_{t}$ is that the former allows an agent in state $s$ to obtain a correct, complete count $x_{s^{\prime}}\geq 0$ of all the agents in state $s^{\prime}$, for any state $s^{\prime}$ that is visible to $s$ (note the count may be zero). Unique IDs are thus not required as there is no risk of counting the same agent twice when receiving communicated counts: either all agents in $s^{\prime}$ have been counted, or no count has yet been obtained for $s^{\prime}$. This simplifies the algorithm and helps preserve agent privacy.
Secondly, uncounted agents cannot be in states for which a count has already been obtained, since the count is complete and correct, even if the count is $x_{s^{\prime}}=0$. Therefore after the $C_{e}$ communication rounds, the $uncounted\_agents$ proportion needs to be uniformly distributed only across the positions in the vector that still have the $\emptyset$ marker, and not across all states as in the general setting. This makes the estimation more accurate in this special setting.
5 Theoretical results
To demonstrate the benefits of the networked architecture by comparison, we also consider (theoretically here and experimentally in Sec. 6) the results of modified versions of our algorithm for centralised and independent learners. In the centralised setting, the Q-network updates of arbitrary agent $i=1$ are automatically pushed to all other agents, and the true global mean-field distribution is always used in place of the local estimate i.e. $\tilde{\hat{\mu}}^{i}_{t}$ = ${\hat{\mu}}_{t}$. In the independent case, there are no links in $\mathcal{G}^{comm}_{t}$ or $\mathcal{G}^{vis}_{t}$, i.e. $\mathcal{E}^{comm}_{t}=\mathcal{E}^{vis}_{t}=\emptyset$.
Networked agents often learn faster than centralised ones in our experiments; we justify theoretically this possibly counterintuitive result here. We first make two strong assumptions that give conditions under which networked agents do outperform centralised ones. The fact that these strong assumptions do not always hold in reality explains why networked agents may not always outperform centralised ones.
Recall that at each iteration $k$ of Alg. 1, after independently updating their policies in Line 14, the population has the policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$. There is randomness in these independent policy updates, stemming from the random sampling of each agent’s independently collected buffer. In Lines 15-20, agents approximate the infinite discounted returns {$V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$ (Def. 3) of their updated policies by computing $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$: the $E$-step discounted return with respect to the empirical mean field generated when agents follow policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$ (i.e. they do not at this stage all follow a single identical policy). We can characterise the approximation as $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$ = $\widehat{V}(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$.
Assumption 1.
Assume that $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$ are sufficiently good approximations so as to respect the ordering of the true values {$V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$, i.e. $\forall i,j\in\{1,\dots,N\}$: $\sigma^{i}_{k+1}>\sigma^{j}_{k+1}\iff V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))>V(\pi^{j}_{k+1},I(\pi^{j}_{k+1})).$
Assumption 2.
Assume that after the $C_{p}$ rounds in Lines 21-27 in which agents exchange and adopt policies from neighbours, the population is left with a single policy such that $\forall i,j\in\{1,\dots,N\}$ $\pi^{i}_{k+1}=\pi^{j}_{k+1}$. [^2] Most simply we can think of Ass. 2 holding if 1) $\tau^{comm}_{k}\rightarrow 0\;\forall k$ such that the softmax essentially becomes a max function, and 2) the communication network $\mathcal{G}^{comm}_{t}$ is static and connected during the $C_{p}$ communication rounds, where $C_{p}$ is larger than the network diameter $d_{\mathcal{G}^{comm}_{t}}$ Under these conditions, previous results on max-consensus algorithms show that all agents in the network will converge on the highest $\sigma^{max}_{k+1}$ value (and hence the unique associated $\pi^{max}_{k+1}$ within a number of rounds equal to the diameter $d_{\mathcal{G}^{comm}_{t}}$ Nejad et al. (2009); Benjamin and Abate (2023). However, policy consensus as in Ass. 2 might be achieved even outside of these conditions, including if the network is dynamic and not connected at every step, given appropriate values for $C_{p}$ and $\tau^{comm}_{k+1}\in\mathbb{R}_{>0}$.
Call the network consensus policy $\pi^{\mathrm{net}}_{k+1}$, and its associated finitely approximated return $\sigma^{\mathrm{net}}_{k+1}$. Recall that the centralised case is where the Q-network update of arbitrary agent $i=1$ is automatically pushed to all the others instead of the policy evaluation and exchange in Lines 15-27; this is equivalent to a networked case where policy consensus is reached on a random one of the policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$. Call this policy arbitrarily given to the whole population $\pi^{\mathrm{cent}}_{k+1}$, and its associated finitely approximated return $\sigma^{\mathrm{cent}}_{k+1}$.
Theorem 1.
Given Ass. 1 and 2, $\mathbb{E}[V(\pi^{\mathrm{net}}_{k+1},I(\pi^{\mathrm{net}}_{k+1}))]>\mathbb{E}[V(\pi^{\mathrm{cent}}_{k+1},I(\pi^{\mathrm{cent}}_{k+1}))]$. Thus in expectation networked agents will increase their returns faster than centralised ones.
Proof.
Recall that before the communication rounds in Line 21 (Alg. 1), the randomly updated policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$ have associated approximated returns $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$. Denote the mean and maximum of this set $\sigma^{\mathrm{mean}}_{k+1}$ and $\sigma^{\mathrm{max}}_{k+1}$ respectively. Since $\pi^{\mathrm{cent}}_{k+1}$ is chosen arbitrarily from $\pi^{i}_{k+1}$}${}_{i=1}^{N}$, it will obey $\mathbb{E}[\sigma^{\mathrm{cent}}_{k+1}]=\sigma^{\mathrm{mean}}_{k+1}$ $\forall k$, though there will be high variance. Conversely, the softmax adoption probability (Line 24, Alg. 1) for the networked case means by definition that policies with higher $\sigma^{i}_{k+1}$ are more likely to be adopted at each communication round. Thus the $\pi^{\mathrm{net}}_{k+1}$ that gets adopted by the whole networked population will obey $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]>\sigma^{\mathrm{mean}}_{k+1}$ (if $\tau^{comm}_{k+1}\rightarrow 0$, it will obey $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]=\sigma^{\mathrm{max}}_{k+1}$ $\forall k$). As such, $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]>\mathbb{E}[\sigma^{\mathrm{cent}}_{k+1}]$, which by Ass. 1 implies the result. ∎
The adoption scheme in Line 24 biases the spread of policies towards those estimated to be better, which, given sufficiently good approximations (Ass. 1), results in higher discounted returns in practice. By choosing updates in a more principled way, networked agents learn faster than the centralised case that adopts updates regardless of quality. This intuition applies even if we loosen Ass. 2 that the networked population converges on a single consensus policy within the $C_{p}$ communication rounds. Similar logic can also be applied to understand why networked agents outperform entirely independent ones, combined with the fact that divergence between policies in the independent case worsens sample complexity over the networked and centralised cases by biasing approximations of the Q-function Yardim et al. (2023); Benjamin and Abate (2023).
Significantly, the communication scheme not only allows us to avoid the undesirable assumption of a central learner, but even to outperform it. Moreover, the benefit of networked communication over centralised learning is even greater in our function approximation setting than in the tabular case of Benjamin and Abate (2023), perhaps due to greater variance in the quality of Q-function estimates in our case.
6 Experiments
We provide two sets of experiments. The first set showcases that our function-approximation algorithm (Alg. 1) can scale to large state spaces for population-independent policies, and that in such settings networked, communicating agents can outperform purely-independent and even centralised agents, and do so by an even greater margin than in the tabular settings from Benjamin and Abate (2023). The second set (given in Sup. Mat. C.4) demonstrates that Alg. 1 can handle population-dependent policies, as well as the ability of Alg. 3 to practically estimate the mean-field distribution locally.
For the types of game used in our demonstrations we follow the gold standard in prior MFG works, i.e. grid-world environments where agents can move in the four cardinal directions or remain in place Laurière (2021); Laurière et al. (2022b); Zaman et al. (2023); Algumaei et al. (2023); Cui et al. (2023); Benjamin and Abate (2023); Wu et al. (2024). We present results from five tasks defined by the agents’ reward/transition functions, all but one of which are coordination tasks. The first two tasks are those used with population-independent policies in Benjamin and Abate (2023), but while they show results for an 8x8 and a ‘larger’ 16x16 grid, our results are for 100x100 and 50x50 grids:
  • Cluster. Agents are rewarded for gathering but given no indication where to do so, agreeing it over time.
  • Target agreement. Agents are rewarded for visiting any of a given number of targets, but the reward is proportional to the number of other agents co-located at the target. Agents must coordinate on which single target they will all meet at to maximise their individual rewards.
See Sup. Mat. C.1 for a full technical description of these tasks, as well as descriptions of additional more complex tasks. In these spatial environments, both the communication network $\mathcal{G}^{comm}_{t}$ and the visibility graph $\mathcal{G}^{vis}_{t}$ are determined by the physical distance from agent $i$; we show plots for various radii, expressed as fractions of the maximum possible distance (the grid’s diagonal length). We discuss hyperparameters in Sup. Mat. C.3.
Figure 1: ‘Target agreement’, pop.-independent, 100x100 grid.
Figure 2: ‘Cluster’, pop.-independent, 100x100 grid.
We evaluate our experiments via two metrics. Exploitability is the most common metric in works on MFGs, and is a measure of proximity to the MFNE. It quantifies how much a best-responding agent can benefit by deviating from the set of policies that generate the current mean-field distribution, with a decreasing exploitability meaning the population is closer to the MFNE. However, there are several issues with this metric in our setting, particularly for our coordination tasks where competitive agents benefit from aligning behaviours, such that it may give limited or noisy information (discussed further in Sup. Mat. C.2.1). We thus also give a second metric, as in Benjamin and Abate (2023): the population’s average discounted return. This allows us to compare how quickly agents are learning to increase their returns, even when ‘exploitability’ gives us limited ability to distinguish between the desirability of the MFNEs to which populations converge.
6.1 Results and discussion
In Figs. 1 and 2, networked agents all significantly outperform both centralised and independent agents in term of average return, despite the centralised agents appearing to have similar exploitability, and the independent agents having similar or slightly lower exploitability. This is because independent agents (and also the centralised ones in Fig. 1) hardly improve their policies at all, such that there is little a deviating agent can do to increase its return in these coordination games, meaning exploitability appears low, despite this being an undesirable equilibrium (see Sup. Mat. C.2 for further discussion on the limited information provided by the exploitation metric). The fact that the networked agents nevertheless significantly outperform the other architectures in terms of average return indicates that communication helps agents to find substantially ‘preferable’ equilibria. Moreover, the margin by which the networked agents can outperform the centralised agents is much greater than in Benjamin and Abate (2023), showing that the benefits of the communication scheme are even greater in non-tabular settings. See Sup. Mat. C.4 for further experiments with large state spaces, as well as others demonstrating the ability of our algorithm to handle two more complex tasks, using population-dependent policies and estimated mean-field observations.
7 Conclusion
We novelly contributed function approximation to the online-learning setting for empirical MFGs, and also contributed two novel algorithms for locally estimating the empirical mean field for population-dependent policies. We have justified theoretically why our networked communication algorithm is able to learn faster than both centralised and independent agents in this function approximation setting, and demonstrated empirically the ability of our algorithms to handle large state spaces and estimate the mean field. Limitations and ways to extend our algorithms are in Sup. Mat. E.
Supplementary material to ‘Networked Communication for Mean-Field Games with Function Approximation and Empirical Mean-Field Estimation’
@@ 7 Conclusion
Appendix A Mean-field estimation algorithm for visibility-based environments
Algorithm 3 Mean-field estimation and communication for environments with $\mathcal{G}^{vis}_{t}$
@@ Appendix A Mean-field estimation algorithm for visibility-based environments
13:  return  mean-field estimates $\{\tilde{\hat{\mu}}^{i}_{t}\}_{i=1}^{N}$
Appendix B Additional remarks on mean-field estimation algorithms
In our Algs. 2 and 3, agents share their local counts with neighbours on the communication network $\mathcal{G}^{comm}_{t}$, and only after the $C_{e}$ communication rounds do they complete their estimated distribution by distributing the uncounted agents along their vectors. An alternative would be for each agent to immediately form a local estimate from their local count obtained via $\mathcal{G}^{obs}_{t}$ or $\mathcal{G}^{vis}_{t}$, which is only then communicated and updated via the communication network. However, we take the former approach to avoid poor local estimations spreading through the network and leading to widespread inaccuracies. Information that is certain (the count) is spread as widely as possible, before being locally converted into an estimate of the total mean field. The same would be the case in our extension proposed in Sec. E for averaging noisy counts, i.e. only the counts would be averaged, with the estimates completed by distributing the remaining agents after the $C_{e}$ communication rounds.
Appendix C Experiments